Answer five questions given the following data (all values are in billions of dollars) Consumption Private Government Exports Imports Expenditures Investment P Investment Purchases Year (GDP) (C x) 2006 13,1789,207 2,220 $ 2,508 s 1481 2.238 2007 13,808 9,710 $ 2.130 2,675 S 1662 S 2.370 2008 14.291 10.035 2087 2878 S 1848 S 2,557 2009 13,939 9,866 1547 S 2918 1583 1975 2010 14,524 10.246 1,795 $ 3.000 1.840 2.357 2011 15,518 10,689 2240 S 3,169 $ 2.106 2,686 2012 16,163 11083 S 2.479 3169 2.194 S 2.763 2013 16,768 11484 2648 $ 3,143 S 2.262 $ 2,770 2014 17.393 11,863 2887 s 3,152 S 2375 $ 2884 2015 18,037 $ 12.284 $ 3057 3,218 2.264 1786 2016 18,569 12.758 30363277 2232 S 2.734 Prey 7 of9 Score answer> SAMSUNG
ent Use the following data to answer the questions below Billions of Dollars Category 500 26 18 45 Consumption Depreciation Retained earnings Gross investment Imports Exports Net foreign factor income 65 16 75 Government purchases Instructions: Enter your responses rounded to the nearest whole number a. How much is GDP? billion b How much is net investment? How much i national income? SAMSUNG
Answer #1

(B) Net Investment, I= Investment - Depreciation= 45-26 = 19

(A) GDP= C + I (from above) + G + (X-M) = 500+ 19+ 75 + (65-55) = 604

(c) National Income= GDP + Net foreign Factor Income = 604+ 16 = 620.